Market Size and Market Share Variances for Small Business Dianes Designs is a small business run out of its owners house. FOperating income $ 35 S56 $ 55 S 90 Page 706 Total Units 80 Sales Actual Results Wolcom. Sign Birdhous. 45 $675 $420 270 $150Help with 3 and 4, please.
Answer #1

Note : As specifically asked Help with 3 and 4, answered for the same only

3.

Values as provided in the issue

Budgeted for month Actual for Month
Product Welcome Sign Bird House Total Welcome Sign Bird House total
units 50 25 75 45 35 80
Sales 1000 250 1250 675 420 1095
V/c -890 -120 -1010 -580 -270 -850
Contribution 110 130 240 95 150 245
F/c -75 -75 -150 -75 -75 -150
Profit 35 55 90 20 75 95
Market 500 200 700 500 175 675

Calculation of market share

Budgeted market share Actual Market Share
Product Diane's Design Market Share Product Diane's Design Market Share
A B C D=B/C A B C D=B/C
Welcome Sign 50 500 10.00% Welcome Sign 45 500 9.00%
Bird House 25 200 12.50% Bird House 35 175 20.00%
Weighted Average Budgeted Contribution Margin Per Unit :
= Budgeted contribution margin per unit for both products/ Budgeted sales per unit for both products
= 240/75
= 3.2

Market size Variance

Market size variance =[(Actual Industry Sales in Units – Budgeted Industry Sales in Units) x Budgeted Market Share Percentage] x Budgeted Average Unit Contribution Margin.
Welcome Sign =[(500-500) x 10.00%] x 3.2
= 0
Bird House = [(175-200) x 12.50%] x 3.2
=10 unfavorable

4. Possible reason for the above variance

Diane’s market size for Welcome Signs remain unchanged while total market for Bird House decreased from budgeted 200 units to actual 175 units . While Diane’s market share for Welcome Signs decreased while it did very well for Birdhouses.

Yet Diane able to sold more units (i.e 80-75 = 5 units).Hence possible reasons for the same may includes changes in quality, price, and seasonal variations etc..